# arXiv:2602.23189v2[math.AP] 02 Mar 2026

### v2

@@ 2.3 Second part of the entropy
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3, then there exits $C>0$ depending on $\|\nabla R_{i}\|_{L^{3}}$ such that
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)-\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)&\leqslant C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}\\ &\quad+C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}.\end{split}$ (2.10)
@@
We have
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i}^{\varepsilon})(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x$ $=\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}^{\varepsilon}+u^{\varepsilon}\cdot\nabla\ln R_{i}^{\varepsilon})\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s,$
for a.a. $t\in(0,T)$ and $i=+,$-. Therefore
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i}^{\varepsilon})(t,x)\,\mathrm{d}x$ $-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x-\Big(\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x\Big)$
$+\int_{0}^{t}\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}u^{\varepsilon}(s,x)\,\mathrm{d}x\,\mathrm{d}s=0.$ (2.11)
We multiply now the mass equation of $R_{\pm}$ by $R_{\pm}^{\varepsilon}/R_{\pm}$ and we can write, for $i=+,$-,
$\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x$ $=\int_{0}^{t}\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}\Big(\partial_{t}R_{i}+u\cdot\nabla R_{i})\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.$
Therefore
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x&=\int_{0}^{t}\int_{\Omega}R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}+u^{\varepsilon}\cdot\nabla\ln_{+}R_{i})(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad-\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}u\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.\end{split}$
@@
Then we use the mass equation of $R_{i}^{\varepsilon}$ in the weak sense taking $\phi=\ln R_{i}$, we obtain
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x&=\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i})(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}(x)\,\mathrm{d}x\\ &\quad-\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}u\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.\end{split}$ (2.12)
Adding (2.11) and (2.12) we get
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x+\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}(u^{\varepsilon}-u)\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s=0.$
Finally, we obtain
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x=-\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
$\qquad-\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u^{\varepsilon}-u)\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.$
@@
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})(u^{\varepsilon}-u)\nabla\ln_{+}R_{\pm}\rvert$ $\leqslant\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\lVert\nabla\ln_{+}R_{\pm}\rVert_{L^{3}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)}.$
We use Poincaré inequality to obtain
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})(u^{\varepsilon}-u)\nabla\ln_{+}R_{\pm}\rvert$ $\leqslant C\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}.$
And similarly we get for the second term
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})\operatorname{div}(u^{\varepsilon}-u)\rvert$ $\leqslant\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}.$
and we reach the desired conclusion. ∎
Corollary 2.6.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3, then there exits $C>0$ depending on $\|\nabla R_{i}\|_{L^{3}}$ such that
${\mathcal{E}}_{2}(t)-{\mathcal{E}}_{2}(0)\leqslant C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\Big(\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}+\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}\Big)+C.$
Proof.
First, note that we have
$f\ln_{+}f=f\ln f-f\ln f1_{f\in(0,1)}$
Using the fact that $x\ln x1_{(0,1)}\leqslant C$, we obtain the following inequality (see [20] for similar inequality)
$f\ln_{+}f\leqslant f\ln f+C.$
It implies that
$R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)\leqslant R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)+CR_{i}$
and then
$\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)$ $-\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)$
$\leqslant$ $\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)-\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)$
$+\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)-\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)+C\int R_{i}(t)$
$\leqslant C\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}}\Big(\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}+\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}\Big)+C\int R_{i}(0)$
where we have used Remark 1.8 and (1.6).
3 Key estimates for the convergence analysis
In this section, we will always denote by $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ a weak solution of System (LABEL:system_eps) in the sense of Definition 2.1 and $(\rho_{+},\rho_{-},\alpha_{+},u)$ a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3.
@@ 3 Key estimates for the convergence analysis
Let us define, for $i=+,$-, $B_{i}^{1}=\{\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert\leqslant 1\}$ and $B_{i}^{2}=\{\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert\geqslant 1\}$. Therefore, we have, for $i=+,$-
$\int_{B_{i}^{1}}\alpha_{i}^{\varepsilon}\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert^{2}\leqslant C\int_{B_{i}^{1}}\alpha_{i}^{\varepsilon}H(\rho_{i}^{\varepsilon}\rvert\rho_{i})\leqslant C\varepsilon^{2}{\mathcal{E}}_{1},$ (3.1)

### v3

@@ 3.2 Logarithmic entropies
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition 2.1, then for all $(\rho_{+},\rho_{-},\alpha_{+},u)$ functions verifying the regularity properties of the limit solution, see Theorem 2.3, we have
$\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)-\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)=$ $\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}$
$+\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u-u^{\varepsilon})$
$-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+\operatorname{div}(R_{i}u)),$
for $a.$ $t\in(0,T)$ and $i=+,$-.
@@
We have for $a.$ $t\in(0,T)$ and $i=+,$-
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i}^{\varepsilon})(t,\cdot)-\int_{\Omega}R_{i,0}^{\varepsilon}\ln R_{i,0}^{\varepsilon}$ $=\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}^{\varepsilon}+u^{\varepsilon}\cdot\nabla\ln R_{i}^{\varepsilon})$
$=\int_{0}^{t}\!\!\!\int_{\Omega}\partial_{t}R_{i}^{\varepsilon}+u^{\varepsilon}\cdot\nabla R_{i}^{\varepsilon}$
$=-\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}u^{\varepsilon}$
where we have used the conservation of the total mass, see (2.3).
@@
Next, we use the mass equation of $R_{i}^{\varepsilon}$ in the weak sense taking $\phi=\ln R_{i}$, we obtain
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i})(t,\cdot)-\int_{\Omega}R_{i,0}^{\varepsilon}\ln R_{i,0}$ $=\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}+u^{\varepsilon}\cdot\nabla\ln R_{i})$
$=\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+u^{\varepsilon}\cdot\nabla R_{i})$
$=\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}(u^{\varepsilon}-u)\nabla\ln R_{i}+\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+u\cdot\nabla R_{i}).$
Subtracting these two equalities, we get
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x$ $-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}(u-u^{\varepsilon})\nabla\ln R_{i}-\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}u^{\varepsilon}-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+u\cdot\nabla R_{i})$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}+\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}(u-u^{\varepsilon})\nabla\ln R_{i}$
$+\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}(u-u^{\varepsilon})-\int_{0}^{t}\!\!\!\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}u-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+u\cdot\nabla R_{i})$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}+\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u-u^{\varepsilon})$
$-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}(\partial_{t}R_{i}+\operatorname{div}(R_{i}u)).$
Proposition 3.5.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition 2.1, and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a solution of System (1.4), verifying the regularity properties of the limit solution, see Theorem 2.3, then there exits $C>0$ depending on $\|\nabla R_{i}\|_{L^{3}}$ such that
${\mathcal{E}}_{2}^{\varepsilon}(t)-{\mathcal{E}}_{2}^{\varepsilon}(0)\leqslant C\int_{0}^{t}\sum_{i=+,-}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)},$
for $a.$ $t\in(0,T)$.
Proof.
Considering $R_{\pm}$ a solution of System (1.4), we can rewrite the expression obtained in Proposition 3.4
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x$ $-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}+\int_{0}^{t}\!\!\!\int_{\Omega}(R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u-u^{\varepsilon}).$
@@
$\int_{\Omega}\lvert(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}\rvert$ $\leqslant\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\lVert\nabla\ln R_{i}\rVert_{L^{3}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)},$
hence, by Sobolev embedding, we obtain
$\int_{\Omega}\lvert(R_{i}^{\varepsilon}-R_{i})(u-u^{\varepsilon})\nabla\ln R_{i}\rvert$ $\leqslant C\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}.$
By Poincaré inequality, we also get for the second term
$\int_{\Omega}\lvert(R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u^{\varepsilon}-u)\rvert\leqslant C\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}.$
We reach the desired conclusion by the conservation of the total mass for $\int R_{i}^{\varepsilon}$ and $\int R_{i}$. ∎
3.3 Control of the densities and the partial densities by the relative entropies
We work under the hypotheses of the main theorem 1.3, in particular, we will always denote by $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ a weak solution of System (LABEL:system_eps) in the sense of Definition 2.1 and $(\rho_{+},\rho_{-},\alpha_{+},u)$ a solution of System (1.4), verifying the regularity properties of Theorem 2.3.
@@ 3.3 Control of the densities and the partial densities by the relative entropies
Choosing $K=[\min(\rho_{-},\rho_{+}),\max(\rho_{-},\rho_{+})]$, we deduce from our assumption $\gamma_{\pm}\geqslant 2$ that, for $\rho_{i}=C_{0}^{1/\gamma_{i}}$ as in the limit system (1.4), we have for a.e $(t,x)\in(0,T)\times\Omega$ and $i=+,$-
$\alpha_{i}^{\varepsilon}|\rho_{i}^{\varepsilon}-\rho_{i}|^{2}+\alpha_{i}^{\varepsilon}|\rho_{i}^{\varepsilon}-\rho_{i}|^{\gamma_{i}}\leqslant C\alpha_{i}^{\varepsilon}H_{i}(\rho_{i}^{\varepsilon}|\rho_{i}),$ (3.1)
where $C$ depends only on $C_{0}$ and $\gamma_{\pm}$.
We begin by establishing the following control of the densities by ${\mathcal{E}}_{1}^{\varepsilon}$.
Proposition 3.6.
If $(\int R_{\pm,0}^{\varepsilon})_{\varepsilon}$ is bounded, then for all $i=+,$- and $p<\gamma_{i}$, there exists $\beta_{1,i}(p),\beta_{2,i}(p)>0$ depending only on $p$ and $\gamma_{\pm}$, and $C_{p}>0$ independent of $\varepsilon$ such that for all $t\geqslant 0$
$\|(\rho_{i}^{\varepsilon}-\rho_{i})(t,\cdot)\|_{L^{p}(\Omega)}\leqslant C_{p}\Big(\varepsilon^{2}{\mathcal{E}}_{1}^{\varepsilon}(t)\Big)^{\beta_{1,i}(p)}+C_{p}\Big(\varepsilon^{2}{\mathcal{E}}_{1}^{\varepsilon}(t)\Big)^{\beta_{2,i}(p)}.$
In the case where $\gamma_{i}=2$ then the convergence holds for $p\leqslant\gamma_{i}$.
Proof.
As all the problem is symmetric in terms of $\gamma_{+}$, $\gamma_{-}$, we can assume without loss of generality that $\gamma_{+}>\gamma_{-}\geqslant 2$. Alternatively, we can exchange $\alpha_{\pm},\rho_{\pm}$ by $\alpha_{\mp},\rho_{\mp}$ in this subsection. Hence $\gamma=\gamma_{+}/\gamma_{-}>1$. The equality of pressures gives
$\rho_{-}^{\varepsilon}=\rho_{+}^{\varepsilon\gamma},$
so that we can write